# MEDebiaser: A Human-AI Feedback System for Mitigating Bias in Multi-label Medical Image Classification

### v2

@@ 4.1.1. Dataset
To evaluate the system, we conducted tests on several datasets, with ChestX-ray14 (Wang et al., 2017) being one of the most representative examples. This dataset includes $112,120$ frontal view chest X-ray images, each at a resolution of $1024\times 1024$, collected from $30,805$ patients between $1992$ and $2015$. These images have been automatically labeled into $14$ different thoracic pathology categories based on radiological diagnostic reports, with each image potentially containing multiple labels. As shown in Figure 9, Appendix B, the label distribution is highly imbalanced, with the most frequent label appearing nearly $20,000$ times and the least frequent only $227$ times. Additionally, there is a noticeable co-occurrence among the labels, as shown in Table 11, Appendix B.
@@ 4.1.3. Interaction
Initially, users can upload their datasets through the Panel View by clicking [Uncaptioned image] (Figure 3- 1). Once the dataset is uploaded, users can observe the proportion and distribution of each label in the Label View via a table (Figure 3- 2). The sorting feature allows users to easily identify labels with smaller proportions, which may be more prone to biases and errors during model training and thus require special attention (DG5). Each label is color-coded in the table, and in the accompanying chord diagram (Figure 3- 3), which is well-suited for showing data with associative relationships within the dataset (Wang et al., 2024a). Each arc connects to its co-occurring labels, with the band thickness representing the frequency of co-occurrence. This diagram enables users to clearly observe label co-occurrence and identify features that may need additional focus (DG5). Next, users can select a model from a predefined set using [Uncaptioned image] (Figure 3- 4). To ensure the model can run on various machines, we provide commonly used CNN models for multi-label scenarios with default settings (batch size = $4$, epoch = $30$), which simplifies the technical setup and allows physicians to proceed with confidence without deep ML knowledge. These default parameters are highly versatile and compatible with different hardware environments, computational capabilities, and workflows in healthcare facilities. For physicians who need to modify parameters or are familiar with ML, clicking on [Uncaptioned image] allows direct adjustment of these parameters and models using [Uncaptioned image] (Figure 3- 5). This flexibility offers more diverse options to optimize training according to their specific needs, training time, and hardware requirements. The user then clicks [Uncaptioned image] (Figure 3- 6), and the preliminary training begins. Upon completion, users can review the performance of each label in the Performance View. Labels that require further attention can be selected by clicking [Uncaptioned image] in the table or by clicking on the arcs in the chord diagram, which will then highlight them in the Attention View.
@@ 4.2.3. Ranking
    (1) inverse $\text{inverse frequency}(c,j)=\frac{1}{M_{cj}+0.01}$
    The small constant $0.01$ ensures stability by preventing division by zero. The overall dependency score for the target label $c$ is then obtained by normalizing these inverse frequencies and summing over the relevant labels. To prevent rarely co-occurring labels from inflating the scores, we include only “positive labels”—those whose co-occurrence with $c$ exceeds a defined threshold—ensuring that weak or incidental associations have no impact on the final score. The dependency score is calculated as (2):
@@
    A dependency score close to $1$ indicates a high likelihood of misclassification due to strong interdependencies, while a score close to 00 suggests lower interdependency and thus fewer challenges in classification.
@@ 4.2.4. Interaction
For the first objective, the Attention View provides a clear display of the Grad-CAM visualization of the model’s decision-making process on the validation set. Users can zoom in on any image by clicking on it, with icons on the top left indicating the accuracy of the model’s label recognition— [Uncaptioned image] for correct predictions and [Uncaptioned image] for incorrect ones. This allows physicians to assess whether the local explanations for the selected labels are reasonable, particularly for symptoms that are less frequent or often co-occur with others (DG2). If an explanation is deemed unreasonable, users can click [Uncaptioned image] (Figure 3- 7) in the bottom right, which will bring the corresponding image into the Modification View. For those reasonably predicted, users can click [Uncaptioned image] (Figure 3- 8) located in the top right corner. This action will save the Grad-CAM of the selected image as polygons, which will then be used in the next round of fine-tuning.
For the second objective, we employ annotation methods familiar to physicians (Hanbury et al., 2017), allowing them to mark areas of interest directly on the image using polygons, which is an intuitive process that does not require ML knowledge, thereby lowering the technical barrier for them to contribute their domain expertise (DG5). In the Modification View, users can enter annotation mode by clicking [Uncaptioned image] (Figure 3- 9). They can then draw polygons by clicking on the image and double-clicking to finish the annotation (Figure 3- 10). Annotating at the pixel level enables the model to focus on more detailed features (DG6). After completing the annotation, users can save their work by clicking [Uncaptioned image] (Figure 3- 11), storing the polygon for the next round of fine-tuning. Annotated images will display [Uncaptioned image] (Figure 3- 12) in the bottom left corner of the images in Attention View. Users can also differentiate between various labels within the same image by selecting from [Uncaptioned image] (Figure 3- 13) and annotating different pixel-level areas accordingly.
In the Recommendation Area (Figure 2-
D
2
), we prioritize images requiring attention by calculating label prediction accuracy, heatmap concentration, and co-occurrence matrix dependency. Users can choose one of these three methods to sort and focus on the images most in need of modification (Figure 3- 14). Additionally, in the Note Area, users can take notes on the image. Compared to the traditional practice of annotating entire datasets, these designs significantly reduce the workload of physicians.
After labeling the biased images and saving the polygon information, users can click [Uncaptioned image] (Figure 3- 15) to start a new round of fine-tuning (DG7). Once completed, the results will be updated in the Performance View.
@@ 4.3. Evaluating Model Performance
In the Performance View, users begin by selecting the label they wish to review from [Uncaptioned image] (Figure 3- 16). The Prediction Results (Figure 2-
E
1
) displays, for each round of fine-tuning, the number of correctly and incorrectly predicted labels on the test set. This visualization enables users to track the trend in correct predictions over time, providing insight into model performance progression. By observing the trends in these charts, physicians can assess whether their modifications are steering the model in a positive direction (DG3). Clicking on the [Uncaptioned image] button reveals the evaluation metrics for each round, including precision, recall (or sensitivity), F1 score, and AUC. The rationale for selecting these specific metrics—while omitting others—is detailed in Table 12Appendix D. To accommodate the varying levels of technical expertise among users, the system adopts a layered presentation strategy. While these parameters are not directly shown in the default view, they remain accessible through interaction for users with machine learning knowledge, such as engineers, to verify that the model remains functional and reliable (DG4). At the same time, physicians are provided with visual trends that reflect performance changes in a more intuitive and accessible manner, aligned with their diagnostic thinking patterns. This design intentionally abstracts away complex terminology, reducing cognitive load for non-expert users, while preserving access to detailed metrics when needed. In doing so, the system strikes a balance between interpretability and transparency, fostering trust and effective collaboration between domain experts and technical practitioners.
In the Comparison Area (Figure 2-
E
2
), test set images are displayed vertically, with the first column showing the Grad-CAM generated after the initial training, and each subsequent column presenting the Grad-CAM after each round of fine-tuning. Images with a green border indicate correct label predictions, while a red border indicates incorrect predictions. By observing these changes in local explanations, physicians can more intuitively understand how the focus of the model’s attention evolves, gaining deeper insights into the model’s decision-making process (DG3).
@@ 5.1.1. Setup and Details
Both experiments were conducted using the ChestX-ray14 dataset and carried out on an NVIDIA 4070 Ti GPU. To reduce the annotation workload for physicians, we selected approximately one-tenth of the original dataset while preserving the label distribution and co-occurrence patterns. The data was split into an 8:1:1 ratio for training, validation, and testing. A DenseNet model pretrained on ImageNet was employed for both experiments, with fine-tuning performed using a learning rate of $41e-$ 4, a batch size of $4$, and $30$ training epochs. To maintain consistency and robustness, we applied standard data augmentation techniques across both experiments.
@@ 5.1.2. Experiment I: With and Without Attention
Procedure. We began by pretraining a DenseNet model on ImageNet, followed by initial training on the ChestX-ray14 training set. The model generated chest X-ray images with Grad-CAM heatmaps and corresponding predictions for both the validation and test sets. Physicians were then tasked with reviewing these images, selecting $100$ examples specifically focusing on the label Pleural_Thickening due to its rarity and higher error rate in prior predictions. Each selected image was meticulously annotated, with polygons drawn around regions of interest, covering all relevant labels.
Using these $100$ annotated images, we employed two model fine-tuning techniques. The first method combined prediction loss with attention loss, as outlined in Section 4.2.2, leveraging Grad-CAM to refine the model’s attention towards the correct regions. The second method relied solely on prediction loss, without any attention-based adjustments.
@@
Result & Analysis. The results, as shown in Table 3, indicate that the model fine-tuned using the combined prediction loss and attention loss approach demonstrated notable improvements in handling the label Pleural_Thickening compared to using prediction loss alone. Specifically, the combined approach achieved a precision of $0.160$, representing a $14.3\%$ relative increase from $0.140$ under prediction loss only. This improvement is also reflected in recall and F1 score. In terms of overall performance, AUC also showed a meaningful enhancement, improving from $0.613$ with prediction loss only to $0.675$ when incorporating attention loss. This suggests a stronger ability to distinguish between positive and negative cases. These quantitative gains underscore the impact of attention modification, which proved particularly beneficial for underrepresented, tail-end labels. By incorporating attention loss, the model was able to more effectively focus on the relevant image regions associated with such rare labels, reducing false negatives, and addressing the inherent challenges posed by imbalanced multi-label data.
@@ 5.1.3. Experiment II: Breadth and Depth
Procedure. We considered two dimensions for Experiment II: breadth and depth. Breadth refers to annotating a larger number of images, and depth focuses on consistently annotating images for a single label. The experimental setup closely mirrored that of Experiment I, utilizing four annotation strategies: 1) annotating $50$ images with Pleural_Thickening; 2) annotating $100$ images with Pleural_Thickening; 3) randomly selecting and annotating $50$ images, and 4) randomly selecting and annotating $100$ images. Efforts were made to maintain consistency in the total number of labels covered across these strategies. For each, the model was fine-tuned using a combined approach of prediction and attention loss.
@@
Metric Preliminary Training 50 Images (Focused) 100 Images (Focused) 50 Images (Random) 100 Images (Random)
@@
Result & Analysis. As shown in Table 4, the model’s performance varied depending on the annotation strategy. Strategies involving a higher number of annotated images consistently yielded in higher precision and AUC compared to those using only $50$ images. Notably, focusing on Pleural_Thickening, which is underrepresented, led to significant improvements in recall. These findings highlight that both annotation depth and breadth play a crucial role in enhancing the model’s generalization capabilities and reducing bias, particularly for tail-end labels.
@@ 5.2.1. Participants
We collaborated with otolaryngology experts from [Blinded for Review], and computer science experts from a local university. We recruited 12 participants, as detailed in Table 5 (physicians: mean age = $40$, SD = $6.56$; engineers: mean age = $32.7$, SD = $4.19$). The group included $6$ otolaryngology experts and $6$ ML experts, with $3$ holding a Master’s degree, $5$ holding an M.D., and $4$ holding a Ph.D.. Notably, seven of the participants had prior experience collaborating with counterparts in medical machine learning projects. Three physicians have no background or limited experience with AI. Participants were randomly paired into $6$ teams consisting of one physician and one engineer.
@@
ID Gender Age Experience1 Degree ID Gender Age Experience2 Degree
@@
  • 1
    Experience stands for Experience with engineers & AI.
  • 2
    Experience stands for Experience with physicians.
@@ 5.2.2. Dataset and Model
In the user study, we utilized an ear endoscopy dataset provided by participants of the study, sourced from [Blinded for Review]. This dataset was collected from $1,272$ patients between $2018$ and $2022$, comprising a total of $3,252$ images, each with a corresponding diagnostic report. We extracted information from the reports and labeled the associated images. Initially, $12$ labels were present in the dataset, but after excluding labels with insufficient representation, we refined it to include $7$ labels across $3,084$ images, referred to as EarEndo. This dataset exhibits significant issues of Imbalanced Distribution and Label Co-occurrence, as shown in Figure 11, Appendix E and  Table 13, Appendix E. In accordance with the hospital’s confidentiality policies, patient names, IDs, addresses, and other personal information were anonymized. Due to the limited dataset size and time constraints of the user study, we instructed the physicians to use DenseNet with the default parameter settings (batch size = $4$, epoch = $30$) throughout the process. The study was conducted on an NVIDIA 4070 Ti GPU.
@@ 5.2.3. Procedure and Task
Figure 4 illustrates the procedure of our user study. Before the study, participants signed privacy and confidentiality agreements and completed a pre-task questionnaire to collect demographic information and task-related background. One author then conducted a 10-minute tutorial on MEDebiaser usage and assigned the tasks for the study (Table 6). Then, each team used MEDebiaser to mitigate bias related to a specific label. During this period, only physicians operated, while physicians and engineers were restricted from communication except during the Evaluating Model Performance stage. The physicians’ interactions with MEDebiaser were screen recorded, and conversations between physicians and engineers were recorded. Afterwards, participants completed a post-task questionnaire individually. Finally, each team participated in an approximately 20-minute interview. The entire user study lasted about one hour on average per participant, who received a \$$10$ token of appreciation upon completion.

### v3

@@ 4.1.1. Dataset
To evaluate the system, we conducted tests on several datasets, with ChestX-ray14 (Wang et al., 2017) being one of the most representative examples. This dataset includes 112,120112,120112 , 120 frontal view chest X-ray images, each at a resolution of $\times 1024$ 1024, collected from 30,80530,80530 , 805 patients between 199219921992 and 201520152015. These images have been automatically labeled into 141414 different thoracic pathology categories based on radiological diagnostic reports, with each image potentially containing multiple labels. As shown in Figure 9, Appendix B, the label distribution is highly imbalanced, with the most frequent label appearing nearly 20,00020,00020 , 000 times and the least frequent only 227227227 times. Additionally, there is a noticeable co-occurrence among the labels, as shown in Table 11, Appendix B.
@@ 4.1.3. Interaction
Initially, users can upload their datasets through the Panel View by clicking [Uncaptioned image] (Figure 3-[Uncaptioned image]). Once the dataset is uploaded, users can observe the proportion and distribution of each label in the Label View via a table (Figure 3-[Uncaptioned image]). The sorting feature allows users to easily identify labels with smaller proportions, which may be more prone to biases and errors during model training and thus require special attention (DG5). Each label is color-coded in the table, and in the accompanying chord diagram (Figure 3-[Uncaptioned image]), which is well-suited for showing data with associative relationships within the dataset (Wang et al., 2024a). Each arc connects to its co-occurring labels, with the band thickness representing the frequency of co-occurrence. This diagram enables users to clearly observe label co-occurrence and identify features that may need additional focus (DG5). Next, users can select a model from a predefined set using [Uncaptioned image] (Figure 3-[Uncaptioned image]). To ensure the model can run on various machines, we provide commonly used CNN models for multi-label scenarios with default settings (batch size = 444, epoch = 303030), which simplifies the technical setup and allows physicians to proceed with confidence without deep ML knowledge. These default parameters are highly versatile and compatible with different hardware environments, computational capabilities, and workflows in healthcare facilities. For physicians who need to modify parameters or are familiar with ML, clicking on [Uncaptioned image] allows direct adjustment of these parameters and models using [Uncaptioned image] (Figure 3-[Uncaptioned image]). This flexibility offers more diverse options to optimize training according to their specific needs, training time, and hardware requirements. The user then clicks [Uncaptioned image] (Figure 3-[Uncaptioned image]), and the preliminary training begins. Upon completion, users can review the performance of each label in the Performance View. Labels that require further attention can be selected by clicking [Uncaptioned image] in the table or by clicking on the arcs in the chord diagram, which will then highlight them in the Attention View.
@@ 4.2.3. Ranking
    (1) $\text{inverse frequency}(c,j)=\frac{1}{M_{cj}+0.01}$
    The small constant 0.010.010.01 ensures stability by preventing division by zero. The overall dependency score for the target label $c$ is then obtained by normalizing these inverse frequencies and summing over the relevant labels. To prevent rarely co-occurring labels from inflating the scores, we include only “positive labels”—those whose co-occurrence with $c$ exceeds a defined threshold—ensuring that weak or incidental associations have no impact on the final score. The dependency score is calculated as (2):
@@
    A dependency score close to 111 indicates a high likelihood of misclassification due to strong interdependencies, while a score close to 0 suggests lower interdependency and thus fewer challenges in classification.
@@ 4.2.4. Interaction
For the first objective, the Attention View provides a clear display of the Grad-CAM visualization of the model’s decision-making process on the validation set. Users can zoom in on any image by clicking on it, with icons on the top left indicating the accuracy of the model’s label recognition— [Uncaptioned image] for correct predictions and [Uncaptioned image] for incorrect ones. This allows physicians to assess whether the local explanations for the selected labels are reasonable, particularly for symptoms that are less frequent or often co-occur with others (DG2). If an explanation is deemed unreasonable, users can click [Uncaptioned image] (Figure 3-[Uncaptioned image]) in the bottom right, which will bring the corresponding image into the Modification View. For those reasonably predicted, users can click [Uncaptioned image] (Figure 3-[Uncaptioned image]) located in the top right corner. This action will save the Grad-CAM of the selected image as polygons, which will then be used in the next round of fine-tuning.
For the second objective, we employ annotation methods familiar to physicians (Hanbury et al., 2017), allowing them to mark areas of interest directly on the image using polygons, which is an intuitive process that does not require ML knowledge, thereby lowering the technical barrier for them to contribute their domain expertise (DG5). In the Modification View, users can enter annotation mode by clicking [Uncaptioned image] (Figure 3-[Uncaptioned image]). They can then draw polygons by clicking on the image and double-clicking to finish the annotation (Figure 3-[Uncaptioned image]). Annotating at the pixel level enables the model to focus on more detailed features (DG6). After completing the annotation, users can save their work by clicking [Uncaptioned image] (Figure 3-[Uncaptioned image]), storing the polygon for the next round of fine-tuning. Annotated images will display [Uncaptioned image] (Figure 3-[Uncaptioned image]) in the bottom left corner of the images in Attention View. Users can also differentiate between various labels within the same image by selecting from [Uncaptioned image] (Figure 3-[Uncaptioned image]) and annotating different pixel-level areas accordingly.
In the Recommendation Area (Figure 2-[Uncaptioned image]), we prioritize images requiring attention by calculating label prediction accuracy, heatmap concentration, and co-occurrence matrix dependency. Users can choose one of these three methods to sort and focus on the images most in need of modification (Figure 3-[Uncaptioned image]). Additionally, in the Note Area (Figure 2-[Uncaptioned image]), users can take notes on the image. Compared to the traditional practice of annotating entire datasets, these designs significantly reduce the workload of physicians.
After labeling the biased images and saving the polygon information, users can click [Uncaptioned image] (Figure 3-[Uncaptioned image]) to start a new round of fine-tuning (DG7). Once completed, the results will be updated in the Performance View.
@@ 4.3. Evaluating Model Performance
In the Performance View, users begin by selecting the label they wish to review from [Uncaptioned image] (Figure 3-[Uncaptioned image]). The Prediction Results (Figure 2-[Uncaptioned image]) displays, for each round of fine-tuning, the number of correctly and incorrectly predicted labels on the test set. This visualization enables users to track the trend in correct predictions over time, providing insight into model performance progression. By observing the trends in these charts, physicians can assess whether their modifications are steering the model in a positive direction (DG3). Clicking on the [Uncaptioned image] button reveals the evaluation metrics for each round, including precision, recall (or sensitivity), F1 score, and AUC. The rationale for selecting these specific metrics—while omitting others—is detailed in Table 12Appendix D. To accommodate the varying levels of technical expertise among users, the system adopts a layered presentation strategy. While these parameters are not directly shown in the default view, they remain accessible through interaction for users with machine learning knowledge, such as engineers, to verify that the model remains functional and reliable (DG4). At the same time, physicians are provided with visual trends that reflect performance changes in a more intuitive and accessible manner, aligned with their diagnostic thinking patterns. This design intentionally abstracts away complex terminology, reducing cognitive load for non-expert users, while preserving access to detailed metrics when needed. In doing so, the system strikes a balance between interpretability and transparency, fostering trust and effective collaboration between domain experts and technical practitioners.
In the Comparison Area (Figure 2-[Uncaptioned image]), test set images are displayed vertically, with the first column showing the Grad-CAM generated after the initial training, and each subsequent column presenting the Grad-CAM after each round of fine-tuning. Images with a green border indicate correct label predictions, while a red border indicates incorrect predictions. By observing these changes in local explanations, physicians can more intuitively understand how the focus of the model’s attention evolves, gaining deeper insights into the model’s decision-making process (DG3).
@@ 5.1.1. Setup and Details
Both experiments were conducted using the ChestX-ray14 dataset and carried out on an NVIDIA 4070 Ti GPU. To reduce the annotation workload for physicians, we selected approximately one-tenth of the original dataset while preserving the label distribution and co-occurrence patterns. The data was split into an 8:1:1 ratio for training, validation, and testing. A DenseNet model pretrained on ImageNet was employed for both experiments, with fine-tuning performed using a learning rate of $1e-4$, a batch size of 444, and 303030 training epochs. To maintain consistency and robustness, we applied standard data augmentation techniques across both experiments.
@@ 5.1.2. Experiment I: With and Without Attention
Procedure. We began by pretraining a DenseNet model on ImageNet, followed by initial training on the ChestX-ray14 training set. The model generated chest X-ray images with Grad-CAM heatmaps and corresponding predictions for both the validation and test sets. Physicians were then tasked with reviewing these images, selecting 100100100 examples specifically focusing on the label Pleural_Thickening due to its rarity and higher error rate in prior predictions. Each selected image was meticulously annotated, with polygons drawn around regions of interest, covering all relevant labels.
Using these 100100100 annotated images, we employed two model fine-tuning techniques. The first method combined prediction loss with attention loss, as outlined in Section 4.2.2, leveraging Grad-CAM to refine the model’s attention towards the correct regions. The second method relied solely on prediction loss, without any attention-based adjustments.
@@
[…]
@@ 5.1.3. Experiment II: Breadth and Depth
Procedure. We considered two dimensions for Experiment II: breadth and depth. Breadth refers to annotating a larger number of images, and depth focuses on consistently annotating images for a single label. The experimental setup closely mirrored that of Experiment I, utilizing four annotation strategies: 1) annotating 505050 images with Pleural_Thickening; 2) annotating 100100100 images with Pleural_Thickening; 3) randomly selecting and annotating 505050 images, and 4) randomly selecting and annotating 100100100 images. Efforts were made to maintain consistency in the total number of labels covered across these strategies. For each, the model was fine-tuned using a combined approach of prediction and attention loss.
@@
Metric Preliminary Training
50 Images
100 Images
50 Images
100 Images
@@ 50 Images
Result & Analysis. As shown in Table 4, the model’s performance varied depending on the annotation strategy. Strategies involving a higher number of annotated images consistently yielded in higher precision and AUC compared to those using only 505050 images. Notably, focusing on Pleural_Thickening, which is underrepresented, led to significant improvements in recall. These findings highlight that both annotation depth and breadth play a crucial role in enhancing the model’s generalization capabilities and reducing bias, particularly for tail-end labels.
@@ 5.2.1. Participants
We collaborated with otolaryngology experts from [Blinded for Review], and computer science experts from a local university. We recruited 12 participants, as detailed in Table 5 (physicians: mean age = 404040, SD = 6.566.566.56; engineers: mean age = 32.732.732.7, SD = 4.194.194.19). The group included 666 otolaryngology experts and 666 ML experts, with 333 holding a Master’s degree, 555 holding an M.D., and 444 holding a Ph.D.. Notably, seven of the participants had prior experience collaborating with counterparts in medical machine learning projects. Three physicians have no background or limited experience with AI. Participants were randomly paired into 666 teams consisting of one physician and one engineer.
@@
ID Gender Age Experience11footnotemark: 1 Degree ID Gender Age Experience22footnotemark: 2 Degree
@@
1 Experience stands for Experience with engineers & AI.
2 Experience stands for Experience with physicians.
@@ 5.2.2. Dataset and Model
In the user study, we utilized an ear endoscopy dataset provided by participants of the study, sourced from [Blinded for Review]. This dataset was collected from 1,2721,2721 , 272 patients between 201820182018 and 202220222022, comprising a total of 3,2523,2523 , 252 images, each with a corresponding diagnostic report. We extracted information from the reports and labeled the associated images. Initially, 121212 labels were present in the dataset, but after excluding labels with insufficient representation, we refined it to include 777 labels across 3,0843,0843 , 084 images, referred to as EarEndo. This dataset exhibits significant issues of Imbalanced Distribution and Label Co-occurrence, as shown in Figure 11, Appendix E and  Table 13, Appendix E. In accordance with the hospital’s confidentiality policies, patient names, IDs, addresses, and other personal information were anonymized. Due to the limited dataset size and time constraints of the user study, we instructed the physicians to use DenseNet with the default parameter settings (batch size = 444, epoch = 303030) throughout the process. The study was conducted on an NVIDIA 4070 Ti GPU.
@@ 5.2.3. Procedure and Task
Figure 4 illustrates the procedure of our user study. Before the study, participants signed privacy and confidentiality agreements and completed a pre-task questionnaire to collect demographic information and task-related background. One author then conducted a 10-minute tutorial on MEDebiaser usage and assigned the tasks for the study (Table 6). Then, each team used MEDebiaser to mitigate bias related to a specific label. During this period, only physicians operated, while physicians and engineers were restricted from communication except during the Evaluating Model Performance stage. The physicians’ interactions with MEDebiaser were screen recorded, and conversations between physicians and engineers were recorded. Afterwards, participants completed a post-task questionnaire individually. Finally, each team participated in an approximately 20-minute interview. The entire user study lasted about one hour on average per participant, who received a \$101010 token of appreciation upon completion.